A boat travels 20 miles upstream in 5hours. Going downstream, it can travel 50 miles in the same amount of time. Find the speed of the current and the speed of the boat in still water.
Answer by mananth(16949)

A boat travels 20 miles upstream in 5hours. Going downstream, it can travel 50 miles in the same amount of time. Find the speed of the current and the speed of the boat in still water.
...
speed of boat = x mph
speed of current = y mph
..
speed against current = 4 mph
speed with current = x+y mph
...
time downstream = time upstream
..
t=d/r
..
5=50/(x+y)
5(x+y)=50
/5
x+y =10
x-y=4
add the equations
2x=14
/2
x=7
..
speed in still water = 7 mph
speed of current = 3 mph
..
m.ananth@hotmail.ca